# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 7 Conclusions
This paper presents a theory for the mass and energy cascade of dark matter on two different levels. On the system level, the cascade in halo mass space leads to the distribution of haloes with respect to halo mass, i.e., the halo mass function. For individual haloes, the cascade along the radial direction leads to the distribution of particles in haloes, i.e., the halo density profile. On both levels, the mass and energy cascade in dark matter flow establishes a statistically steady state to continuously release the system energy and maximize the system entropy. The key feature of the statistically steady state is scale-independent rates of cascade such that the statistical structures of the haloes are self-similar and scale-free, and there is no net accumulation of mass and energy on any intermediate scales.
@@
For the mass cascade in halo mass space, the net mass transfer is upward in a bottom-up fashion, i.e., an inverse cascade that is consistent with the hierarchical structure formation. Two distinct ranges are identified, i.e., a propagation range with a scale-independent rate of mass transfer $\varepsilon_{m}$ below a characteristic mass $m_{h}^{*}$ and a deposition range to actively consume the mass cascaded from small scales to grow halo larger than $m_{h}^{*}$ (Fig. 3). The inverse mass cascade leads to a random walk of haloes in the mass space with a position-dependent waiting time $\tau_{g}$. The distribution of haloes in halo mass space, i.e., the halo mass function, is naturally given by the solution of the corresponding Fokker-Planck equation for halo random walk (double-$\lambda$ mass function in Eq. (33)). This approach for halo mass function is simple without resorting to a spherical or elliptical collapse model. The non-Gaussian features in the density field developed at small scales are highly expected to be an important signature of the inverse mass cascade. Since haloes have finite kinetic and potential energy, there also exists an inverse cascade of kinetic energy at a rate of $\varepsilon_{u}$ from small to large mass scales and a direct cascade of the potential energy at a rate of $1.4\varepsilon_{u}$ from large to small mass scales (Fig. 8). Universal scaling laws in mass space can be developed with small-scale permanence for halo group mass $m_{g}$ (Fig.5).
The N-body system in an expanding background is not energy conserved. The total energy continuously decreases with time (Fig. 1). This is facilitated by the energy cascade from large to small haloes in halo mass space and from large to small scales in individual haloes, such that the energy is "dissipated" on small scales. The energy cascade establishes a statistically steady state to continuously release halo energy and maximize halo entropy. In these haloes, the outflow of halo mass due to the Hubble expansion balances the inflow due to the halo mass accretion and self-gravity (Fig. 9). The net mass transfer vanishes due to the stable clustering hypothesis. The kinetic energy is inversely transferred from small to large scale $r$ at a $r$-independent rate $\varepsilon(m_{h},z)\propto m_{h}^{2/3}a^{-1}$. In contrast, the potential energy is directly cascaded from large to small scales at a rate of $1.4\varepsilon$ (Fig. 11). The outflow of mass can be described by the random walk of particles with a position-dependent waiting time $\tau_{gr}$. The distribution of particles in haloes, i.e., the halo density profile, can be analytically derived from the corresponding Fokker-Planck equation (double-$\gamma$ profile in Eq. (80)). Universal scaling laws (Eqs. (70)) can be identified with small-scale permanence for the halo density profile (Fig. 13). The different inner density slopes of simulated haloes can be explained by the nonzero net mass flux and energy flux in haloes (Eq. (84)).
@@
Since the waiting time and jumping length for particle random walk in haloes depends on the particle mass $m_{p}$ (Eq. (72)), new constraints on the particle mass can be identified from the physical constraints on these quantities. Based on the assumption that the waiting time should be greater than the Planck time (the smallest unit of time) or the jumping length should be greater than the Planck length (the smallest unit of length), we propose a new constraint for particle mass $m_{p}\geq 10^{-15}$kg or $10^{12}$eV (Eqs. (90), Eq. (93), (100), and (107)). These constraints exclude the standard WIMPs and strongly suggest a heavy dark matter scenario (superheavy right-handed neutrinos, etc.). This constraint is also consistent with the particle mass obtained by extending the established scaling laws to the smallest free-streaming scale (Xu, 2022b).
Acknowledgments
@@
Two datasets underlying this article, that is, halo-based and correlation-based statistics of dark matter flow, are available on Zenodo (Xu, 2022c, d), along with the accompanying presentation slides ’A comparative study of dark matter flow & hydrodynamic turbulence and its applications’ (Xu, 2022e). All data are also available on GitHub (Xu, 2022f).
Appendix A Mass cascade in halo mass space
@@
This section presents some additional details on the inverse mass cascade in halo mass space in matter dominant era.
A.1 Mass redistribution among halo groups
@@
All haloes in the halo subsystem can be grouped into haloes with the same mass $m_{h}$ or particle number $n_{p}$. We focus on the mass transfer (cascade) between halo groups of different sizes. Similarly to the concept of halo merger trees, the starting point is to define functions that describe the mass redistribution among halo groups at different redshifts $z$. The forward mass redistribution function $D_{FM}\left(z_{1},n_{p1},z_{2},n_{p2}\right)$ describes the mass fraction of a halo group of size $n_{p1}$ at the redshift $z_{1}$ that is inherited from the halo group of size $n_{p2}$ at an earlier redshift $z_{2}$. Similarly, the backward mass redistribution function $D_{BM}\left(z_{1},n_{p1},z_{3},n_{p3}\right)$ can be defined as the mass fraction of a halo group of size $n_{p1}$ at redshift $z_{1}$ that will be passed to the halo group of size $n_{p3}$ at a later redshift $z_{3}$. These functions precisely quantify the progenitor and inheritor of the total mass of a given halo group. Figure 16 provides an example for two functions. For given $z_{1}$, $n_{p1}$, and $z_{2}$ or $z_{3}$, the normalization condition requires,
@@
Figures. 17 and 18 plot the forward and backward redistribution functions as a function of the halo group size $n_{p2}$ or $n_{p3}$ for five different group sizes $n_{p1}$=2, 5, 10, 20, 50, respectively. Halo groups of five different sizes $n_{p1}$ inherit and pass their mass to a distribution of halo group sizes. The interaction among halo groups is shown to be local on the mass scale. This demonstrates that smooth and minor merging is dominant over the major merging between two haloes of comparable mass. The halo group of size $n_{p1}$ inherits or passes most of its mass via merging/breaking between haloes of similar (neighboring) size to $n_{p1}$ and single mergers. Therefore, there are two peaks for forward/backward mass redistribution functions at around $n_{pi}\approx n_{p1}$ (halo groups of similar size) and $n_{pi}\approx 1$ (single mergers), where i=2 or 3 for forward or backward functions, respectively.
Groups of large haloes inherit/pass their mass from/to a wider distribution of halo sizes. In comparison, groups of small haloes inherit/pass their mass from/to a relatively narrower distribution of halo sizes. Both mass redistribution functions are not symmetric about the halo size $n_{p1}$, with more mass inherited from halo groups below the size $n_{p1}$ and more mass passed to halo groups above the size $n_{p1}$, i.e., a net transfer of mass to large scales. The sharp peaks for halo groups of smaller size and the widespread distribution for halo groups of larger size in Figs. 17 and 18 indicate that small haloes have relatively longer lifespans and can exist for a longer time such that most small haloes will remain in the same group at a later redshift. Large haloes tend to have a relatively shorter lifespan. Halo lifespan will be further discussed in Section 4.4.
@@
Figure 19: The net mass redistribution function $D_{NM}$ as a function of halo group size $n_{p2}$ for five different halo group sizes $n_{p1}$, with $D_{NM}<0$ indicating that the halo group of size $n_{p1}$ inherits more mass from the halo group of size $n_{p2}$ than the mass it passes to the group of the same size $n_{p2}$. The net effect of mass redistribution is that haloes are transferring mass from smaller scales to larger scales, i.e., an inverse mass cascade in mass space. In contrast, direct mass cascade refers to the mass transfer from larger to smaller scales.
@@
    Asymmetry: mass transfer across halo groups is a two-way process with mass cascading both upward and downward in the halo mass space. However, the net mass transfer is upward, i.e., the structure formation proceeds in a “bottom-up” fashion. The mass redistribution functions of a given halo size $n_{p1}$ are asymmetric about $n_{p1}$, with more mass inherited from halo groups smaller than $n_{p1}$ (via halo merging) and less mass inherited from halo groups larger than $n_{p1}$ (via halo breaking-up). Larger-size haloes accrete their mass from a relatively wider size range of haloes, while smaller-size haloes accrete their mass from a narrower size range of haloes.
@@
A.2 Time scales in mass cascade
In this section, we try to develop a few elementary ideas about the time and mass scales for inverse mass cascades. We observe that there exists a broad spectrum of halo sizes. This can be a direct result of maximizing system entropy (Xu, 2023a). The smallest halo is often created by gravitational collapse at the smallest scale that merges with other haloes and passes their mass onto larger haloes. The larger haloes are themselves transitory and pass their mass to even larger haloes, and so on. At every instant t, there is a continuous cascade of mass from the smallest to the largest mass scales that we assumed to be a characteristic mass $m_{h}^{*}$

### v4

@@ 6 Conclusions
In this paper, we propose that, between the fully linear and fully nonlinear regimes, there exists a universal transition range centered on a characteristic mass $m_{h}^{*}$, within which a quasi-universal spectrum with index nn=-1 is established by gravity and becomes largely independent of the specifics of the initial conditions. We identify and develop the physical mechanism that enables this universality as a cascade of mass and energy across scales, and we derive the attendant universal scaling laws governing both the halo mass function and the halo density profiles.
@@
We formulate the mass and energy cascade at two distinct but coupled levels. Globally, a cascade in halo mass space governs the redistribution of mass among haloes and thereby determines the halo mass function. Locally, within individual haloes, a cascade directed along the radius governs how the particles are arranged, and thus determines the halo density profile. On both levels, the cascade drives the system toward a statistically steady state that continuously releases energy and maximizes entropy. A defining property of this steady state is a scale-independent cascade rate: statistical halo structures are self-similar and scale-free, and there is no net accumulation of mass or energy at intermediate scales.
Because an N-body system in an expanding background has a decreasing total energy (Fig. 17), the global decrease or "dissipation" of energy facilitates the cascades: energy is injected around scale $m_{h}^{*}$, transferred from large to small haloes in mass space and from large to small radii within haloes, and is ultimately dissipated on small scales through both smooth halo merging and the motions of particles in haloes. The cascade establishes a statistically steady state with scale-independent rates. We find that kinetic energy is transferred inversely from small to large scales at a rate $\varepsilon(m_{h},z)\propto m_{h}^{2/3}a^{-1}$, while potential energy undergoes a direct cascade from large to small scales at a rate of $1.4\varepsilon$ (Fig. 14).
@@
For the mass cascade in the halo mass space, the net mass transfer is upward in a bottom-up fashion, i.e., consistent with hierarchical structure formation. Two distinct regimes emerge, i.e., a propagation range with a scale-independent rate of mass transfer $\varepsilon_{m}$ below a characteristic mass $m_{h}^{*}$ and a deposition range to actively consume the mass cascaded from small scales to grow halo larger than $m_{h}^{*}$ (Fig. 7). The inverse mass cascade leads to a random walk of haloes in the mass space with a position-dependent waiting time $\tau_{g}$. The distribution of haloes in the halo mass space, i.e. the halo mass function, is naturally given by the solution of the corresponding Fokker-Planck equation for halo random walk (double-$\lambda$ mass function in Eq. (26)). This approach for the halo mass function is simple without resorting to a spherical or elliptical collapse model. The non-Gaussian features in the density field around scales $m_{h}^{*}$ are highly expected to be an important signature of the inverse mass cascade. Since haloes have finite kinetic and potential energy, there also exists an inverse cascade of kinetic energy at a rate of $\varepsilon_{u}$ from small to large mass scales and a direct cascade of potential energy at a rate of $1.4\varepsilon_{u}$ from large to small mass scales (Fig. 10). Universal scaling laws can be developed with small-scale permanence for halo group mass $m_{g}$ (Fig.9).
For the energy cascade in haloes, the kinetic energy is transferred outward, while the potential energy is transferred inward. The origin of halo structure can be described by the random walk of particles with a position-dependent waiting time $\tau_{gr}$. The distribution of particles in haloes, i.e., the halo density profile, can be analytically derived from the corresponding Fokker-Planck equation (double-$\gamma$ profile in Eq. (61)). Universal scaling laws (Eqs. (71)) can be identified with small-scale permanence for the halo density profile (Fig. 11).
@@
Since the waiting time and jumping length for particle random walk in haloes depend on the particle mass $m_{p}$ (Eq. (86)), new constraints on the particle mass can be identified from the physical constraints on these quantities. Based on the assumption that the waiting time should be greater than the Planck time (the smallest unit of time) or the jumping length should be greater than the Planck length (the smallest unit of length), we propose a new constraint for particle mass $m_{p}\geq 10^{-15}$kg or $10^{12}$GeV (Eqs. (98), Eq. (101), (108), and (115)). These constraints exclude the standard WIMPs and strongly suggest a heavy dark matter scenario (superheavy right-handed neutrinos, etc.). This constraint is also consistent with the particle mass obtained by extending the established scaling laws to the smallest free-streaming scale (Xu:2022-Postulating-dark-matter-partic).
Author contributions
Z.J. Xu performed the conception, data analysis, and writing.
Funding
@@
Two datasets underlying this article, that is, halo-based and correlation-based statistics of dark matter flow, are available on Zenodo (Xu:2022-Dark_matter-flow-dataset-part1; Xu:2022-Dark_matter-flow-dataset-part2), along with the presentation ’A comparative study of dark matter flow & hydrodynamic turbulence and its applications’ (Xu:2022-Dark_matter-flow-and-hydrodynamic-turbulence-presentation). All data are also available on GitHub (Xu:Dark_matter_flow_dataset_2022_all_files).
Appendix A N-body simulations and data
The large-scale cosmological Illustris simulation (Illustris-1-Dark) (NELSON:2015-The-illustris-simulation) was used to demonstrate and validate the concepts. Illustris is a suite of large-volume cosmological dark matter only and hydrodynamical simulations. The selected Illustris-1-Dark is a dark matter only simulation that has a cosmological volume of 106.5Mpc3 and 18203 DM particles for a high mass resolution. Each DM particle has a mass around $m_{p}=7.6\times 10^{6}M_{\odot}$. The gravitational softening length is around 1.4 kpc. The simulation has cosmological parameters of dark matter density $\Omega_{DM}=0.2726$, dark energy density $\Omega_{DE}=0.7274$ at $z=0$, and Hubble constant $h=0.704$. Some key parameters of N-body simulations are listed in Table 2.
For cross-validation, the cosmological N-body simulations carried out by the Virgo consortium were also used. A detailed description can be found in (Frenk:2000-Public-Release-of-N-body-simul; Jenkins:1998-Evolution-of-structure-in-cold). The SCDM simulation of the Virgo consortium focuses on matter-dominant dark matter-only simulations with a standard CDM power spectrum (SCDM). The same set of simulation data has been widely used in several different studies, from clustering statistics (Jenkins:1998-Evolution-of-structure-in-cold) to the formation of halo clusters in large-scale environments (Colberg:1999-Linking-cluster-formation-to-l), and testing models for halo abundance and mass functions (Sheth:2001-Ellipsoidal-collapse-and-an-im). This simulation has a lower mass resolution with particle mass $m_{p}=2.27\times 10^{11}{M_{\odot}/h}$. The simulation box is around 240 Mpc/h, where h is the dimensionless Hubble constant in units of 100 km/s/Mpc.
The friends-of-friends algorithm (FOF) was used to identify all haloes in each simulation that depend only on a dimensionless parameter b, which defines the linking length $b\left({N/V}\right)^{{-1/3}}$, where $V$ is the volume of the simulation box. In this work, haloes were identified with a linking length parameter of $b=0.2$. Identifying all haloes of different sizes, all dark matter particles were divided into halo particles with a total mass $M_{h}$ and out-of-halo particles that do not belong to any halo. Therefore, $M_{h}$ is the total mass of all haloes. We focus on the evolution of mass and energy in haloes of different mass $m_{h}$. All haloes were grouped into halo groups of various sizes according to the halo mass $m_{h}$ (or $n_{p}$, the number of particles in the halo), where $m_{h}=n_{p}m_{p}$. The total mass for a halo group of mass $m_{h}$ is $m_{g}=m_{h}n_{h}$, where $n_{h}$ is the number of haloes in that group. Two different cosmological simulations were used to demonstrate the fundamental concepts. The same approach can be easily extended to other cosmological simulations.
Table 2: Virgo (SCDM) and Illustris (Illustris-1-Dark) parameters
Run $\Omega_{0}$ $h$ L (Mpc/h) N $m_{p}$ $M_{\odot}/h$ $l_{soft}$ (Kpc/h)
[…]
Appendix B Cosmic energy evolution
In this section, we analyze the temporal evolution of the global energy in a self-gravitating, collisionless dark-matter flow (SG-CFD). Although the global energy, obtained by averaging across all scales, does not resolve inter-scale energy transfer itself, its systematic decline over time provides the driving force necessary to initiate and sustain the cascade on a global scale. The corresponding scale-to-scale transfer of mass and energy is presented in Section 3.3 and is fundamentally distinct from the Press–Schechter formulations.
The equations of motion for N collisionless particles in comoving coordinates $\boldsymbol{\mathrm{x}}$ and physical time t read (Peebles:1980-The-Large-Scale-Structure-of-t):
where N particles have equal mass $m_{p}$. The Hubble parameter $H\left(t\right)={\dot{a}/a}$. Here, $H$ has a "damping" effect, which leads to a decrease in total energy of the N-body system (Eq. (121) and Fig. 17). In the matter-dominant era, the Hubble parameter satisfies $Ht=2/3$.
Next, we will derive the energy evolution based on the equations of motion (Eq. (116)). We first introduce a transformed time variable s as ${ds/dt}=a^{p}$, where p is an arbitrary exponent. In terms of the new time variable $s$, the original Eq. (116) can be transformed to
where $\boldsymbol{\mathrm{F}}_{i}$ is the resultant force on particle i of all other particles, while $P_{s}$ is the total specific potential energy in the comoving coordinates. Equation (117) reduces to the original Eq. (116) when $p=0$. With p=2p=-2, the first-order derivative vanishes in Eq. (117) and s is the time variable for integration in N-body simulations. By setting p=1p=-1, s is the conformal time. Setting p=3/2p=-{3/2} along with $H_{0}^{2}=H^{2}a^{3}$ for the matter-dominant era, the equation of motion becomes
In this transformed equation for the matter era, the scale factor a is not explicitly involved. The time-dependent Hubble parameter $H$ is replaced by the Hubble constant $H_{0}$. This transformation offers significant convenience for the evolution of cosmic energy (Xu:2022-The-evolution-of-energy--momen).
We first identify the transformation between the velocity $\boldsymbol{\mathrm{v}}_{i}$ in the time variable $s$ and the peculiar velocity $\boldsymbol{\mathrm{u}}_{i}$,
The energy evolution of the N-body system can be obtained by multiplying $\boldsymbol{\mathrm{v}}_{i}=d\boldsymbol{\mathrm{x}}_{i}/ds$} on both sides of Eq. (117) and adding the equation of motion for all particles together (Xu:2022-The-evolution-of-energy--momen). An exact and simple equation (in time variable $s$) for the specific kinetic energy $K_{s}$ and the potential energy $P_{s}$ can be obtained as
$\frac{dK_{s}}{ds}+2HK_{s}(p+2)a^{-p}+a^{-(3+2p)}\frac{dP_{s}}{ds}=0.$ (120)
Setting $p=0$ and using the relations in Eq. (119), the exact cosmic energy equation for the energy evolution of the N-body system reads as
$\frac{\partial E_{y}}{\partial t}+H\left(2K_{p}+P_{y}\right)=0,$ (121)
which describes the energy evolution in an expanding background. Here $K_{p}$ is the peculiar kinetic energy, $P_{y}$ is the potential energy in physical coordinates, and $E_{y}=K_{p}+P_{y}$ is the total specific energy. This is also known as the Layzer-Irvine equation in the literature (Irvine:1961-Local-Irregularities-in-a-Univ; Layzer:1963-A-Preface-to-Cosmogony--I--The).
In the matter-dominant era with HtHt=2/3, the cosmic energy evolution (Eq. (121)) admits a linear solution
where $\varepsilon_{u}<0$ is an important physical constant (unit: m2/s3), the rate of the energy cascade. It represents the rate of time variation of cosmic energy. The solution in Eq. (122) can be directly validated by N-body simulations. Figure 17 presents the evolution of the energy from the Virgo simulation (SCDM). The (specific) kinetic energy $K_{p}$ and potential energy $P_{y}$ were calculated as the mean energy of all dark matter particles in the N-body system. Figure 17 confirms the solution in Eq. (122), i.e., a linear increase of $K_{p}$ with time and a negative potential energy $P_{y}=-1.4K_{p}$. Two key messages are:
i) According to the virial theorem, in virial equilibrium, $2K_{p}=n_{e}P_{y}$, where $n_{e}$ is an exponent of the interaction potential $r^{n_{e}}$. For standard gravity, the exponent $n_{e}=-1$. Since gravity is the only interaction in the N-body system, we would expect $n_{e}=-1$ for an N-body system. However, the virial equilibrium can never be reached by the non-equilibrium N-body system (e.g., SG-CFD of dark matter flow). Instead of the virial equilibrium, a statistically steady state is established to continuously release system energy and maximize system entropy, which is manifested by the mass and energy cascade. In that particular state, the solution of Eq. (122) is valid such that an "effective" exponent $n_{e}=-10/7$ is obtained, which deviates from -1;
ii) The total energy $E_{y}=-0.4K_{p}$ decreases with time. In analogy to hydrodynamic turbulence, the total energy of the N-body system in an expanding background decreases with time, as if "dissipated" due to the expanding background, even though there is no viscous force in collisionless dark matter. Here, we admit that cosmic energy decreases with time and estimate the rate of energy "dissipation":
where $u_{0}\equiv u\left(t=t_{0}\right)\approx 350$km/s is the one-dimensional velocity dispersion of all dark matter particles, and $t_{0}\approx$13.7 billion years is the physical time at the present epoch or the age of the universe. Different simulations may have slightly different values of $u_{0}$ due to different cosmological parameters. However, the rate of energy "dissipation" $\varepsilon_{u}\sim-10^{-7}m^{2}/s^{3}$ should be a good estimate.
Hereafter, the quoted word "dissipation" stands for the decrease in total energy caused by the expanding background, in contrast to the actual dissipation caused by the viscous force in hydrodynamic turbulence. In the statistically steady state, the energy "dissipated" on small scales should balance the energy cascaded from large scales. Therefore, Eq. (122) also suggests an inverse cascade of kinetic energy from small to large scales at a rate of $\varepsilon_{u}$ and a direct cascade of potential energy at a rate of $1.4\varepsilon_{u}$. The total energy is directly cascaded to small scales at a rate of $0.4\varepsilon_{u}$ to provide the energy that is "dissipated" on small scales. In Sections 3 and 4, we illustrate how the mass and energy cascades are initiated by the global-scale energy "dissipation" and proceed in halo mass space via halo merging and in individual haloes via the random walk of particles, and their impacts on structure formation and evolution.
Appendix C More details on the mass cascade
@@
This section presents some additional details on the inverse mass cascade in halo mass space in the matter-dominant era.
C.1 Mass redistribution among halo groups
@@
All haloes in the halo subsystem can be grouped into haloes with the same mass $m_{h}$ or particle number $n_{p}$. We focus on mass transfer (cascade) between halo groups of different sizes. Similarly to the concept of halo merger trees, the starting point is to define functions that describe the mass redistribution among halo groups at different redshifts $z$. The forward mass redistribution function $D_{FM}\left(z_{1},n_{p1},z_{2},n_{p2}\right)$ describes the mass fraction of a halo group of size $n_{p1}$ at the redshift $z_{1}$ that is inherited from the halo group of size $n_{p2}$ at an earlier redshift $z_{2}$. Similarly, the backward mass redistribution function $D_{BM}\left(z_{1},n_{p1},z_{3},n_{p3}\right)$ can be defined as the mass fraction of a halo group of size $n_{p1}$ at redshift $z_{1}$ that will be passed to the halo group of size $n_{p3}$ at a later redshift $z_{3}$. These functions precisely quantify the progenitor and inheritor of the total mass of a given halo group. Figure 18 provides an example for two functions. For given $z_{1}$, $n_{p1}$, and $z_{2}$ or $z_{3}$, the normalization condition requires,
@@
Figures. 19 and 20 plot the forward and backward redistribution functions as a function of the halo group size $n_{p2}$ or $n_{p3}$ for five different group sizes $n_{p1}$=2, 5, 10, 20, 50, respectively. Halo groups of five different sizes $n_{p1}$ inherit and pass their mass to a distribution of halo group sizes. The interaction among halo groups is shown to be local on the mass scale. This demonstrates that smooth and minor merging is dominant over the major merging between two haloes of comparable mass. The halo group of size $n_{p1}$ inherits or passes most of its mass via merging/breaking between haloes of similar (neighboring) size to $n_{p1}$ and single mergers. Therefore, there are two peaks for forward/backward mass redistribution functions at around $n_{pi}\approx n_{p1}$ (halo groups of similar size) and $n_{pi}\approx 1$ (single mergers), where i=2 or 3 for forward or backward functions, respectively.
Groups of large haloes inherit/pass their mass from/to a wider distribution of halo sizes. In comparison, groups of small haloes inherit/pass their mass from/to a relatively narrower distribution of halo sizes. Both mass redistribution functions are not symmetric about the halo size $n_{p1}$, with more mass inherited from halo groups below the size $n_{p1}$ and more mass passed to halo groups above the size $n_{p1}$, i.e., a net transfer of mass to large scales. The sharp peaks for halo groups of smaller size and the widespread distribution for halo groups of larger size in Figs. 19 and 20 indicate that small haloes have relatively longer lifespans and can exist for a longer time such that most small haloes will remain in the same group at a later redshift. Large haloes tend to have a relatively shorter lifespan. Halo lifespan will be further discussed in Section 3.1.
@@
    Asymmetry: mass transfer across halo groups is a two-way process with mass cascading both upward and downward in the halo mass space. However, the net mass transfer is upward, i.e., the structure formation proceeds in a “bottom-up” fashion. The mass redistribution functions of a given halo size $n_{p1}$ are asymmetric about $n_{p1}$, with more mass inherited from halo groups smaller than $n_{p1}$ (via halo merging) and less mass inherited from halo groups larger than $n_{p1}$ (via halo breaking-up). Larger-size haloes accrete their mass from a relatively wider range of haloes, while smaller-size haloes accrete their mass from a narrower size range of haloes.
@@
Figure 21: The net mass redistribution function $D_{NM}$ as a function of halo group size $n_{p2}$ for five different halo group sizes $n_{p1}$, with $D_{NM}<0$ indicating that the halo group of size $n_{p1}$ inherits more mass from the halo group of size $n_{p2}$ than the mass it passes to the group of the same size $n_{p2}$. The net effect of mass redistribution is that haloes are transferring mass from smaller scales to larger scales, i.e., an inverse mass cascade in mass space, in contrast ot a direct mass cascade.
C.2 Time scales in mass cascade
In this section, we try to develop a few elementary ideas about the time and mass scales for inverse mass cascades. We observe that there exists a broad spectrum of halo sizes. This can be a direct result of maximizing system entropy (Xu:2023-Maximum-entropy-distributions-of-dark-matter). The smallest halo is often created by gravitational collapse at the smallest scale that merges with other haloes and passes their mass onto larger haloes. The larger haloes are themselves transitory and pass their mass to even larger haloes, and so on. At every instant t, there is a continuous cascade of mass from the smallest to the largest mass scales that we assumed to be a characteristic mass $m_{h}^{*}$